OpenStudy (anonymous):

why is tan(90+Theta)=-cot Theta?

OpenStudy (shubhamsrg):

you agree that tan(90-x) = cot x?

OpenStudy (anonymous):

Yep

OpenStudy (shubhamsrg):

just substitute x as -x over there..

OpenStudy (anonymous):

so it means that since tan(90-x)=cot x, tan(90-(-x))=-cotx?

OpenStudy (shubhamsrg):

cot(-x) = -cotx ..thats why..

OpenStudy (anonymous):

ok thx!:)